.

April 4, 2025 Eric Babson babson@math.ucdavis.edu 1094 Cragmont Avenue Berkeley CA, 94708

Dear Mr President:

I am concerned about privacy.

Untrained staffers now have access to detailed personal information about tens of millions of us. Data is from the Labor, Education and Health Departments, the Social Security Administration and the Office of Personnel Management with information on student loans, medicare payments, employment, grants, contracts and unaccompanied minors. In several cases the same person is known to have access to many of these but their identities and intentions remain largely hidden while various law suits allege dangers from cyberattacks to retaliatory firing. Lax corporate protections have caused most data breaches though hackers got twenty million federal records a decade ago and Russia infiltrated during your last term. President Biden tightened things up and people have felt safe enough to file taxes, get passports and generally work with the federal government. This confidence is now eroding.

Please assure me that you will protect data.

My pride in the US derives in part from public trust.

Sincerely, Eric Babson